# Decoding Generalization from Memorization in Deep Neural Networks

### v1

@@ 3.1 Experimental Setup
Once the model is trained, we apply MASC on each layer of the network with respect to different subspaces. For MLP models, all the MASC experiments were performed for all the layers in the network including on the input (after it is pre-processed). For CNN models and AlexNet models, the experiments were performed on flatten layer (Flat) and fully connected layers (FC). While we ran the experiments on the input layer for CNNs, we did not do so for AlexNet.
3.2 Terminology
The general terminology used in this work is as follows:
Model Training Accuracy: The model accuracy on the training set with corrupted labels.
Model Testing Accuracy: The model accuracy on the testing data set with true labels.
Minimum Angle Subspace Classifier (MASC) Accuracy on Corrupted Training: Training accuracy of MASC on training data set with corrupted labels was used in determining the subspaces.
Minimum Angle Subspace Classifier (MASC) Accuracy on Original Training: Training accuracy of MASC on training data with respect to true training labels.
Minimum Angle Subspace Classifier (MASC) Accuracy on Testing: Testing accuracy of MASC on testing data set with true labels was used.
4 Enhanced generalization ability in memorized models
@@
Figure 1: Minimum Angle Subspace Classifier (MASC) accuracy over the layers of the network when the data is projected onto corrupted training subspaces with the indicated corruption degree, for multiple models/datasets. Rows corresponds to plots with the same corruption degree & the columns correspond to the models, as noted. Training accuracy (dashed line) & testing accuracy (dotted line) of the model is shown. FC corresponds to fully connected layer with $ReLU$ activation whereas Flat corresponds to flatten layer without $ReLU$ activation. The number of class-wise PCA components of these models are shown in Figure 16 in section A.9 of the Appendix.
@@
Models trained with corrupted labels have high training accuracy (on corrupted labels) while also having low testing accuracy (Zhang et al., 2021). We ask if we can decode the representations of the hidden layers of these memorized models to obtain better generalization.
@@
To do so, we build a Minimum Angle Subspace Classifier (MASC) using class-conditioned corrupted training subspaces obtained from the memorized models’ hidden layer outputs. MASC is performed layer-wise for all the layers of the network independently as described in Section 3. MASC accuracy on corrupted training data, MASC accuracy on original training data, and MASC accuracy on testing data over the layer of MLP trained on MNIST, CNN trained on Fashion-MNIST, AlexNet trained on CIFAR-100 are shown in Figure 1. $SGD$ optimizer (Qian, 1999) was used for training MLP models, whereas $Adam$ optimizer (Kingma, 2014) was used for other models.
@@
Importantly, for every corrupted model we have (with non-zero corruption degree), except those with 100% corruption degree, we find that our Minimum Angle Subspace Classifier (MASC) in at least one layer has better testing accuracy than the corresponding model itself. In many cases, the MASC testing accuracy is dramatically better than that of the model. This is remarkable, because, in addition to the layerwise outputs, the MASC used precisely the same information (including the same corrupted training dataset) that was available to the model itself, and yet is able to extract better generalization. This suggests that the model retains significant latent generalization ability, which is not captured in its own test-set performance. In most models, the same MASC, especially on the later layers, also approaches perfect accuracy on the corrupted training set, indicating that this improved generalization happens concurrently with memorization of training data points with shuffled labels. Below, we make more specific observations on the performance of the models.
@@
With generalized models i.e. those with 0% corruption degree, at the later layers of the network, it is observed that in most of the cases MASC accuracy on training data approaches the models training accuracy. Similarly, MASC accuracy on testing data is comparable to or performed better than the models’ test accuracy.
@@
Even for high corruption degrees, we find that the MASC performs well. For example, with 80% corruption degree, which implies that approximately 72% of the training labels have been changed, we observed good MASC testing accuracy in many cases. Notably, the MASC test accuracy on the later layers is over 80% on MLP-MNIST, in comparison to 34% test accuracy by the model. Similarly, MASC test accuracy on one of the layers is about 75% for CNN-Fashion-MNIST, in contrast to 25% model test accuracy.
@@
Not only does the MASC have better accuracy than the model on the test data but it also does well on the training data with the true labels. Although the model has memorized the training data with corrupted labels, outputs from certain layers have the ability to predict the trained true labels. For example, in MLP-MNIST, for low to moderate degrees of corruption, MASC on the middle layer (FC (512)) has good accuracy on the true training labels, while also retaining good accuracy on the test set. With 40% corruption degree, approximately 36% are changed labels and yet the model has good accuracy on the true training labels in at least one layer of the network. e.g. MLP-MNIST has over 90% true training accuracy at layer FC(512), CNN-Fashion-MNIST has approximately 85% in Flat (576) layer & AlexNet-CIFAR-100 has approximately 60% in FC (4096) layer. This means that almost 20% of those labels are predicted correctly even though the model was trained for 500 epochs or has reached high training accuracy on corrupted labels. In the process of doing this, the model does not have any direct information about the true labels and neither does the MASC.
@@
One way to think about a deep network, is as one that successively transforms input representations in a manner that aids in good prediction performance. Therefore, performance of the MASC on the input is a good baseline measure to assess if subsequent layers have favorable accuracies. Naively, for models trained with corrupted data, one would expect layered representations that enable the model to do well on the corrupted training data, but not do well on the test data or the training data that have true labels. While this expectation seems to hold with respect to the model itself, we find that the layer-wise representations do not necessarily follow this expectation. That is, MASC applied to subsequent layers, often have better true training accuracy and test accuracy than the MASC applied to the input, suggesting that the deep network does indeed transform the data in a manner more amenable to correct prediction, even if its labels are dominated by noise.
MLP model trained on CIFAR-10 with $SGD$ optimizer is shown in Section A.10 along with MLP models trained on CIFAR-10 and MNIST with $Adam$ having qualitatively similar results. We ran some preliminary experiments with Dropout as a regularizer. To do so, we have trained CNN on MNIST, Fashion-MNIST, CIFAR-10 and AlexNet on CIFAR-100 with dropout. The details and results are provided in Section A.11.
5 Generalization via true training labels with memorized models
Figure 2: Minimum Angle Subspace Classifier (MASC) accuracy over the layers of the network when the data set is projected onto corrupted subspace and subspace corresponding to true training labels. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models as noted. Training and testing accuracy of the model is shown. FC corresponds to fully connected layer with ReLU activation whereas Flat corresponds to flatten layer without ReLU activation. The respective number of class-wise PCA components for true training label subspaces of the models is shown in Figure 17 in section A.9 of the Appendix.
While the previous section demonstrated improved generalization performance by the MASC, we want to investigate if there exist better subspaces that can offer superior generalization performance. To this end, we consider the setting where the true label identities of the training set are known post training with corrupted labels. Can we extract significantly high training as well as testing performance in this case from the layerwise outputs of the network? To do so, we build MASC using subspaces obtained from training data with true labels. It is a priori unclear if MASCs trained in this manner will have high accuracy. Since the network trained assuming different labels for many of the datapoints, it is conceivable that class-wise subspaces corresponding to true labels lack structure and predictive power. We find, however, that these possibilities do not bear out.
MASC accuracy on original training data and on testing data projected on true training label subspace over the layers of the same networks from Section 4 is shown in Figure 2. For comparison, MASC accuracy on corrupted training data and testing data projected on corrupted training subspace is also shown. We find that, in many cases, accuracies on the true training labels, as well as the test set are dramatically better here than with the experiments where subspaces were determined for the corrupted training data. In fact the MASC test accuracies for the corrupted models (with non-zero corruption degree) are sometimes fairly close to the test accuracy of the uncorrupted model.
Strikingly, even for models trained with 100% corruption degree, in most cases, the MASC retains significant accuracy on the true training labels as well as the test set. This is in spite of the fact that the model itself has chance-level test-set accuracy. For example, MASC classifier has 95% test labels accuracy in last FC(2048) layer for MLP-MNIST, 69% test labels accuracy for Flat(576) layer in CNN-Fashion-MNIST, and 4% test labels accuracy for Flat(256) layer in AlexNet-CIFAR-100.
The results here are proof of principle that suggest the existence of subspaces which allow one to extract significantly high generalization performance on models trained with datapoints whose labels are shuffled to a remarkably high degree. This has two implications. On the one hand, it demonstrates that models trained with very high label noise, surprisingly, retain the latent ability to generalize very well. On the other hand, it suggests that development of new techniques to identify favorable subspaces could help markedly boost generalization performance of models, whose training data is known to have label noise.
Similar results were seen with respect to $Adam$ optimizer over the MLP layers trained on MNIST and CIFAR-10, and MLP trained on CIFAR-10 with $SGD$ optimizer as shown in Figure 21 with its respective class-wise PCA components available in Figure 22 in the Appendix. Results for model with regularization are shown in Figure 29 and 31 with its respective class-wise PCA components available in Figure 30 and 32.
The Appendix also describes a control experiment with MASC accuracies on a random initialization of the network (Section A.6, as well as comparison with early stopping test accuracies (Section A.4). We also have results corresponding to more models and datasets (Section A.8) and details of experiments for AlexNet-Tiny ImageNet (Section A.5).
6 Inducing Memorization in uncorrupted models
@@
Figure 3: Minimum Angle Subspace Classifier (MASC) accuracy over the layers of the generalized network when the data set is projected onto corrupted training subspaces with the indicated corruption degree. Rows corresponds to plots which have the same corruption degree & the columns correspond to the generalized models as noted. Training & testing accuracy of the generalized model is shown. FC corresponds to fully connected layer with $ReLU$ activation whereas Flat corresponds to flatten layer without $ReLU$ activation. The respective number of class-wise PCA components of the models is shown in Figure 18 in Section A.9 of the Appendix.
Conversely, we examined if we could build a MASC classifier on a model trained on true training labels, with the goal of memorizing training data whose labels are corrupted to varying degrees.
To do this, we take generalized models, i.e. models trained with uncorrupted training data. We then shuffle the labels of the training set to some corruption degree and construct the corresponding class-specific subspaces with respect to the layerwise outputs of the model. We then build a MASC classifier corresponding to these subspaces.
MASC accuracy on original training data and MASC accuracy on testing data over the layer same networks from Section 4 are shown in Figure 3. Additional results are available in the Appendix in Figures 23, 33, 35 and 39 and their respective class-wise PCA components are available in Figure 24, 34, 36.
Interestingly, we find that for uncorrupted model with modest model test accuracies (i.e. AlexNet-CIFAR-100), the MASC classifiers described above have high accuracies on the corrupted training set. Conversely, in most uncorrupted models with high model test accuracies (i.e. MLP-MNIST and CNN-Fashion-MNIST), we find that these MASC classifiers have more modest accuracies on the training set with corrupted labels. One exception to this, is in Figure 23 of the Appendix, where we have MLP-Adam-MNIST models with high model test accuracy. Yet, we find that a MASC classifier on the first FC(2048) layer trained with training labels corrupted to 100% corruption degree has over 90% accuracy on training data with corrupted labels. Also, MASC classifiers often have test accuracies that approach or exceed uncorrupted model test accuracies, even though they correspond to corrupted subspaces (see e.g. AlexNet on CIFAR-100).

### v2

@@ 3.3 Minimum Angle Subspace Classifier Algorithm (MASC)
We apply MASC on each layer of the network with respect to different subspaces. For MLP models, all the MASC experiments were performed for all the layers in the network including on the input (after it is pre-processed). For CNN models and AlexNet models, the experiments were performed on flatten layer (Flat) and fully connected layers (FC). For ResNet-18 model, we evaluated nine layers – L0, L0-1, L0-2, L0-3, L1, L2, L3, L4, and the average-pool (avg_pool) layer – containing 16,384; 16,384; 16,384, 16,384; 16,384; 8,192; 4,096; 2,048; and 512 neurons respectively. Here, L0 denotes the layer immediately before the L1 block; L0-1, L0-2, and L0-3 are intermediate outputs within the L1 block; and L1–L4 correspond to the outputs of successive residual blocks. All layer outputs were flattened prior to analysis. While we ran the experiments on the input layer for CNNs, we did not do so for AlexNet or ResNet-18.
3.4 Leveraging MASC to retrain the model
Here, the idea is to use one of the layerwise MASC classifiers in order to relabel the corrupted training set. This relabeled training set is then used to retrain the existing model. To determine the layer whose MASC classifier we will use, we find the layer whose MASC classifier generalizes best. To this end, we first split the test data set into 80%-20%. We use the MASC accuracy on the corrupted subspaces in the validation set (created from 20% of the test dataset) to identify the model’s best-layer. Then, using the best-layer MASC predictions, we relabel the corrupted training dataset. We train with the relabeled corrupted training dataset for upto 30 epochs and perform early stopping with patience of 3 by considering the 20% test dataset as a validation dataset; this validation dataset was not used in reporting test accuracy. A similar process was followed while working with subspaces corresponding to true labels. The test accuracy on the models is calculated with respect to the 80% test dataset, obtained in the aforementioned split. A schematic of the retraining process using MASC is shown in Figure 9.
4 Enhanced innate generalization to true labels in memorized models
@@
Figure 3: Minimum Angle Subspace Classifier (MASC) accuracy over the layers of the network when the data is projected onto corrupted training subspaces with the indicated corruption degree, for multiple models/datasets. Rows corresponds to plots with the same corruption degree & the columns correspond to the models, as noted. Training accuracy (dashed line) & test accuracy (dotted line) of the model is shown. FC corresponds to fully connected layer with ReLUReLU activation whereas Flat corresponds to flatten layer without ReLUReLU activation. The number of class-wise PCA components of these models are shown in Figure 30. SGDSGD optimizer (Qian, 1999) was used for training MLP models, whereas AdamAdam optimizer (Kingma, 2014) was used for other models. ResNet-18 has layer outputs of size 16,384 for L0-L1, followed by 8,192 (L2), 4,096 (L3), 2,048 (L4), and 512 (avg_pool) layer.
@@
Models trained with corrupted labels have high training accuracy (on corrupted labels) while also having low accuracy on the test set with true labels (Zhang et al., 2017; 2021). We ask if we can decode the representations of the hidden layers of these memorized models to obtain better generalization to true labels.
@@
To do so, we build a probe that we call Minimum Angle Subspace Classifier (MASC) using class-conditioned corrupted training subspaces obtained from the memorized models’ hidden layer outputs. MASC is performed layer-wise for the layers of the network independently. More details on MASC are available in the Methods section. MASC accuracy on corrupted training data, MASC accuracy on original training data (with true labels), and MASC accuracy on test data (with true labels) over the layers of MLP trained on MNIST, CNN trained on Fashion-MNIST, AlexNet trained on CIFAR-100, and ResNet-18 trained on CIFAR-10 for various randomly-chosen fractions of label corruption in training data (i.e. corruption degrees) are shown in Figure 3. Likewise, results for MLP trained on CIFAR-10, CNN trained on MNIST & CIFAR-10 and AlexNet trained on Tiny ImageNet are presented in Figure 27.
@@
Importantly, for every corrupted model we have (with non-zero corruption degree), except those with 100% corruption degree, we find that our Minimum Angle Subspace Classifier (MASC) in at least one layer (with one exception [^8] ResNet-18 trained on CIFAR-10 with 20% corruption degree is the lone exception. See Figure 3 and Table 1 for the corresponding results.) has better test accuracy than the corresponding model itself. Table 1 reports by what percentage the MASC classifier outperformed the model for the best such layer, for each model. In Table 6, we also report the accuracy difference between the MASC classifier and the model for the best such layer, for each model. In many cases, the MASC test accuracy is dramatically better than that of the model. This is remarkable, because, in addition to the layerwise outputs, MASC used precisely the same information (including the same corrupted training dataset) that was available to the model itself, and yet is able to extract better generalization to true labels. This suggests that the model retains significant latent generalization to true labels, which is not captured in its own test-set performance. In many models, the same MASC, especially on the later layers, also approaches perfect accuracy on the corrupted training set, indicating that this improved generalization to true labels can happen concurrently with memorization of training data points with shuffled labels. Below, we make more specific observations on the performance of the models.
Table 1: Percentage by which the MASC classifier (run on the best layer) outperformed the model’s test accuracy when the data is projected onto corrupted training subspaces. The best layer corresponds to the one that has the highest measured MASC test accuracy among the layers for the said model/dataset. The accuracies in each case are averaged over three runs and are rounded to the second decimal place. Some of the detailed results are available in Appendix, as indicated.
Corruption degree 20% 40% 60% 80%
MLP-MNIST 10.93% 31.63% 75.04% 159.93%
MLP-CIFAR-10 (Appendix) 9.90% 24.42% 46.97% 64.75%
CNN-MNIST (Appendix) 9.81% 37.03% 98.69% 201.06%
CNN-Fashion-MNIST 7.49% 33.50% 84.93% 189.00%
CNN-CIFAR-10 (Appendix) 2.29% 6.26% 27.03% 60.17%
AlexNet-CIFAR-100 33.58% 53.10% 64.86% 45.00%
AlexNet-Tiny ImageNet (Appendix) 27.50% 53.46% 45.38% 14.16%
ResNet-18-CIFAR-10 -0.41% 20.34% 67.28% 119.16%
@@
With generalized models i.e. those with 0% corruption degree, at the later layers of the network, it is observed that in most of the cases MASC accuracy on training data approaches the models training accuracy. Similarly, MASC accuracy on test dataset is comparable to or performed better than the models’ test accuracy, with the exception of the ResNet-18 model.
Figure 4: MASC test accuracy over the layers of the network when the data is projected onto corrupted training subspaces with the indicated corruption degree. Best model test accuracy corresponds accuracy of the test data of the model if early stopping was used.
@@
Even for high corruption degrees, we find that MASC performs well. For example, with 80% corruption degree, which implies that approximately 72% of the training labels have been changed, we observed good MASC test accuracy in many cases. Notably, the MASC test accuracy on the later layers is over 80% on MLP-MNIST, in comparison to 34% test accuracy by the model. Similarly, MASC test accuracy on one of the layers is about 75% for CNN-Fashion-MNIST, in contrast to 25% model test accuracy. Even for larger models/datasets such as AlexNet-CIFAR-100, MASC test accuracy outperforms the model test accuracy by 45%, for training sets with 80% corruption degree. Likewise, for ResNet-18-CIFAR-10, several layers exhibit MASC test accuracies that exceed the model’s test accuracy.
@@
Not only does MASC have better accuracy than the model on the test data but, when applied to some layers, it also does well on the training data with the true labels. Although the model has memorized the training data with corrupted labels, outputs from certain layers have the ability to predict the trained true labels. For example, in MLP-MNIST, for low to moderate degrees of corruption, MASC on the middle layer (FC (512)) has good accuracy on the true training labels, while also retaining good accuracy on the test set. With 40% corruption degree, approximately 36% are changed labels and yet the model has good accuracy on the true training labels in at least one layer of the network. e.g. MLP-MNIST has over 90% true training accuracy at layer FC(512), CNN-Fashion-MNIST has approximately 85% in Flat (576) layer & AlexNet-CIFAR-100 has approximately 60% in FC (4096) layer. This means that almost 20% of those labels are predicted correctly even though the model was trained for 500 epochs or has reached high training accuracy on corrupted labels. In the process of doing this, the model does not have any direct information about the true labels and neither does MASC.
For a subset of models, we also compare the results of test accuracy of MASC on trained model with early stopping model accuracy. MASC accuracy over the layers of the network when the data is projected onto corrupted training subspaces is shown in Figure 4. Best model test accuracy and trained model test accuracy are shown for reference. Best model test accuracy corresponds to the accuracy of the test data of the model if early stopping was used.
For MASC when the data is projected onto corrupted training subspace, in AlexNet-CIFAR-100, the MASC in at least one layer shows better performance than the best model test accuracy for less than 60% corruption degree. For MLP-MNIST, the best model (early stopping) maintains over 90% accuracy even when the data is corrupted up to 80%. Despite the increase in corruptions (except 100% corruption), the accuracy of the last layer remains close to that of the best model accuracy. For CNN-Fashion-MNIST (except 100% corruption), in at least one layer MASC performance is near to that of best model test accuracy.
@@
One way to think about a deep network, is as one that successively transforms input representations in a manner that aids in good prediction performance. Therefore, performance of MASC on the input is a good baseline measure to assess if subsequent layers have favorable accuracies. Naively, for models trained with corrupted data, one would expect layered representations that enable the model to do well on the corrupted training data, but not do well on the test/training data that have true labels. While this expectation seems to hold with respect to the model itself, we find that the layer-wise representations do not necessarily follow this expectation. That is, MASC applied to subsequent layers, often have better true training accuracy and test accuracy than MASC applied to the input, suggesting that the deep network does indeed transform the data in a manner amenable to better generalization to true labels, even if its labels are dominated by noise.
5 Generalization to true labels comparison: MASC versus other probes
@@
Figure 5: Test accuracies for different probes over the layers of the network. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models as noted. Test accuracy of the model and MASC are shown for comparison. FC corresponds to fully connected layer with ReLU activation whereas Flat corresponds to flatten layer without ReLU activation.
Given that MASC is a probe on layerwise outputs, it is natural to ask how a few other probes might perform in the memorization setting. Accordingly, we have used five different probes [^9] Experimental setup is provided in Section C.1 on the layer of the deep neural networks namely, Logistic Regression (LR) , K-Nearest Neighbor (KNN), Linear Discriminant Analysis (LDA), Quadratic Discriminant Analysis (QDA), and Nearest Class Mean (NCM). Figure 5 reports the test accuracies of all probes across the layers of MLP-MNIST, CNN-Fashion-MNIST, and AlexNet-CIFAR-100. Model and MASC test accuracies are overlaid for comparison. Results for additional models are provided in Section C.2. We also report computational cost (GFLOPS) for different probes over the layer of the network in Section C.3.
We find that, although all probes achieve broadly comparable test accuracies – with no consistent pattern regarding which probe performs best at different corruption levels – their computational costs (GFLOPs) differ substantially. For AlexNet, the probes ranked from highest to lowest computational cost are: QDA, LDA, KNN, MASC, LR, and NCM. For the CNN models, the ordering is KNN, followed by QDA and LDA (which have identical cost), MASC, LR and NCM. For the MLP models, the sequence is KNN, QDA, LDA, MASC, LR, and NCM. Across all models, MASC consistently exhibits lower computational cost compared to the KNN, QDA, and LDA probes, although its cost remains higher than that of LR and NCM.
Given this computational profile, we focus on results comparing the test accuracy of MASC with LR and NCM; the corresponding results are detailed in Section C.4. The three probes display distinct trends across layers and corruption degrees. On AlexNet–CIFAR-100 and MLP–MNIST at 20%–40% corruption, MASC surpasses both NCM and LR on most layers. For CNN-based models, NCM generally performs best, with MASC typically ranking above LR. Notably, MASC matches or exceeds NCM in specific cases such as CNN–MNIST and CNN–Fashion-MNIST at 20%, 40%, and 60% corruption, at the Flat(576) layer.
While our initial focus had been on MASC, it is interesting that other probes also have comparable performance, and indeed, this performance isn’t always correlated with that of MASC. That is, in some cases, these probes perform better than MASC in some layers and worse than MASC in other layers. This suggests that latent representations that contain information useful for generalization to true labels may manifest in different forms, which result in different probes decoding them with differing accuracies. This phenomenon requires deeper investigation, which has been beyond the scope of the present paper.
6 Generalization to true labels via true training label subspaces in memorized models
Figure 6: Minimum Angle Subspace Classifier (MASC) accuracy over the layers of the network when the data set is projected onto corrupted subspace and subspace corresponding to true training labels. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models as noted. Training and test accuracy of the model is shown. FC corresponds to fully connected layer with ReLU activation whereas Flat corresponds to flatten layer without ReLU activation. The respective number of class-wise PCA components for true training label subspaces of the models is shown in Figure 32. SGDSGD optimizer was used for training MLP models, whereas AdamAdam optimizer was used for other models. ResNet-18 has layer outputs of size 16,384 for L0-L1, followed by 8,192 (L2), 4,096 (L3), 2,048 (L4), and 512 (avg_pool) layer.
Figure 7: MASC test accuracy over the layers of the network when the data set is projected onto subspace corresponding to true training labels. Best model test accuracy corresponds accuracy of the test data of the model if early stopping was used.
While Section 4 demonstrated improved generalization to true labels by MASC, here, we investigate if there exist subspaces that can offer even better generalization to true labels. To this end, we consider the setting where the true label identities of the training set are known, after training with corrupted labels is complete. Can we extract significantly high training as well as test performance in this case from the layerwise outputs of the network? To do so, we build MASC using subspaces obtained from training data with true labels. It is a priori unclear if MASCs trained in this manner will have high accuracy. Since the network trained assuming different labels for many of the datapoints, it is conceivable that class-wise subspaces corresponding to true labels lack structure and predictive power. We find, however, that these possibilities do not bear out.
MASC accuracy on original training data and on test data projected on true training label subspace over the layers of the same networks is shown in Figure 6 and Figure 28. For comparison, MASC accuracy on corrupted training data and test data projected on corrupted training subspace is also shown. We find that, in many cases, accuracies on the true training labels, as well as the test set are dramatically better here than with the experiments where subspaces were determined for the corrupted training data. In Table 7, we show by what percentage the MASC classifier outperformed the model for the best layer for corruption degrees 20%, 40%, 60% and 80%. In Table 8, we also report the accuracy difference between the MASC classifier and the model for the best such layer, for each model. In fact, the MASC test accuracies for the corrupted models (with non-zero corruption degree) are sometimes fairly close to the test accuracy of the uncorrupted model.
Notably, even for models trained with 100% corruption degree, in most cases, MASC retains significant accuracy on the true training labels as well as the test set. This is in spite of the fact that the model itself has chance-level test-set accuracy. For example, MASC classifier has 95% test accuracy in the last FC(2048) layer for MLP-MNIST, 69% test accuracy for Flat(576) layer in CNN-Fashion-MNIST, and 4% test accuracy for Flat(256) layer in AlexNet-CIFAR-100.
The results here are proof of principle that suggest the existence of subspaces which allow one to extract significantly high generalization to true labels on models trained with datapoints whose labels are shuffled to a remarkably high degree. This has two implications. On the one hand, it demonstrates that models trained with very high label noise, surprisingly, retain the latent ability to generalize very well. On the other hand, it suggests that development of new techniques to identify favorable subspaces could help markedly boost generalization to true labels of models, whose training data is known to have label noise.
The results comparing MASC test accuracy over the layers of the network when the data is projected onto true training subspaces with early stopping model accuracy is shown in Figure 7.
The Appendix presents MASC comparison results across different PCA variance thresholds, along with an analysis of MASC’s computational and time complexity. The Appendix also describes a control experiment with MASC accuracies on a random initialization of the network. We have results corresponding to MLP trained on CIFAR-10, CNN trained on MNIST and CIFAR-10, and AlexNet trained on Tiny ImageNet in the Appendix for all the experiments. We also have a section comparing MLP models trained on MNIST and CIFAR-10 with SGD and Adam optimizer.
In Section H, we also investigating the latent memorization capabilities of uncorrupted models. Here, conversely, we ask how well a network trained on true labels can manifest memorization of an arbitrary relabeling of its training data. More specifically, we built a MASC classifier on a model trained on true training labels, with the goal of memorizing training data whose labels are corrupted to varying degrees post hoc. Interestingly, we find a dichotomy in model behavior here, with some models trained on specific datasets having the propensity to memorize to a high degree, whereas others not demonstrating such ability.
7 Leveraging MASC to retrain the base memorized model for improved generalization to true labels
Taking into account the better generalization to true labels using MASC on memorized models, in this section, we ask the following question. Can we use the MASC classifier to retrain the existing model to achieve better generalization to true labels?
Details of the pipeline for retraining existing models, leveraging MASC are already presented before. For different corruption degrees, test accuracy [^10] Note that the test accuracy on the models is calculated with respect to the 80% test dataset. before and after retraining with relabeled data using MASC (corrupted and true subspaces) for MLP-MNIST, MLP-CIFAR-10, CNN-MNIST, CNN-Fashion-MNIST, CNN-CIFAR-10 models are shown in Figure 8. Similar results for AlexNet-CIFAR-100, AlexNet-Tiny ImageNet models are shown in Figure 42.
In order to study the dynamics of accuracy during retraining, unencumbered by the early stopping criterion, we also performed a similar experiment without using early stopping, for 10 epochs. The results for model before training, model after retraining on MASC corrupted subspace, and model after retraining on MASC subspace with true labels over the 10 epochs for all model-dataset pairs with various corruption degrees are shown in Figure 46.
In Figure 8 and 42, we find that for some models (MLP-MNIST, CNN-MNIST, CNN-Fashion-MNIST, CNN-CIFAR-10) with non-zero corruption degrees, there is an improvement in the test accuracy of models retrained using relabelling with MASC on corrupted subspaces, in comparison to the models’ test accuracy before retraining (existing models). Indeed, in some cases, the improvement is quite significant, especially for larger corruption degrees (that are below 100% corruption degree).
Figure 8: Test accuracies averaged over three runs on the 80% test dataset is plotted for different model-dataset pairs for various corruption degrees and for various models/MASC classifiers. Model before retraining corresponds to the existing memorized model. Model after retraining on MASC corrupted subspace corresponds to model trained with training dataset relabeled using MASC corrupted subspace predictions on the best layer. Model after retraining on MASC true subspace corresponds to model trained with training dataset relabeled using MASC subspaces corresponding to true label predictions on the best layer. MASC test accuracy on the best layers for corrupted and true label subspaces on existing corrupted models (before retraining) are shown for comparison. The best layer was identified using a validation set that was carved out of the test dataset, for this experiment. Error bar represents the range on three different runs.
However, for some models (MLP-CIFAR-10, AlexNet-CIFAR100), the accuracy gains due to such retraining appear marginal. Indeed, in some cases (MLP-CIFAR10, AlexNet-Tiny ImageNet for 20% corruption degree), there is a decrease in the test accuracy with such retraining. In order to study why, we checked the fraction of incorrect labels in the relabeled training dataset and compared it with the same measure for the existing corrupted training dataset. For the corruption degrees 20%, 40%, 60%, and 80%, these results are plotted in Figure 43. Table 9 lists the exact values of the same.
In particular, it turns out that for MLP-CIFAR-10 and AlexNet-CIFAR-100 this fraction is almost equal to the fraction on the existing corrupted dataset; for MLP-CIFAR-10 the fraction is marginally higher for the relabeled dataset and for AlexNet-CIFAR-100, it is marginally lower. This simply implies that the MASC classifier on the best layer that uses corrupted subspaces does roughly as well on the training set with true labels as the existing model, while surprisingly, the same MASC classifier is able to perform significantly better than the existing model on the test-set with true labels (See light orange bar in Figure 8 and 42). With regard to retraining, it would seem that the relabeled training isn’t more effective than the existing corrupted training set in training the model, which possibly reflects in the lack of significant improvement in test accuracy. More broadly, this suggests that MASC’s better generalization to true labels isn’t necessarily accompanied by better training set performance on true labels. This phenomenon requires a more detailed future investigation.
Secondly, for AlexNet-Tiny ImageNet with 20% corruption degree, the relabeled training set has significantly fewer fraction of correct labels (92.46% lower) than in the existing corrupted training set. We wanted to determine to what extent this lower fraction is driven by previously incorrect label predictions (in the existing corrupted training dataset) being predicted correctly (in the relabeled set), versus previously correct predictions being relabeled incorrectly. For all models, these numbers are visualized in Figures 44 & 45 and Tables 10 & 11 list corresponding numbers. We find for the case of AlexNet-Tiny ImageNet with 20% corruption degree that there is a small fraction (5.94%) of previously incorrect labels that are correctly relabeled and a large fraction (24.46%) of correctly labeled points that are incorrectly labeled by the MASC classifier trained on the corrupted label subspaces. Notably, even though this MASC classifier while doing significantly worse on the training data than the model happens to do markedly better than the model in test accuracy. As before, we think that the poor performance of the retrained model is driven by the fact that relabeling results in a dataset with larger fraction of incorrect labels.
Thirdly, in some cases (e.g., AlexNet-Tiny ImageNet with 40%, 60%, 80% corruption degree), we observe that even though relabeling results in a somewhat larger fraction of incorrect labels, the test accuracy of the retrained model is slightly better. This suggests that retraining performance is not simply driven by fraction of correct labels, but that specifics of which points are relabeled can drive retraining performance in ways that remain to be investigated.
With respect to models retrained on MASC true label subspaces, it was observed that the test accuracy of such models usually performed significantly better than the models before retraining.
By-and-large, we find that the retrained models that use MASC corrupted subspace relabeling don’t have better test performance than the corresponding MASC classifiers. However, it would be interesting to apply MASC on the retrained models to see if that would further improve generalization to true labels performance.